

********************************************************************
UARS WEEKLY STATUS REPORT - 15 JAN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
PEM is now under PMON control! OBC activation tests of the high energy
particle spectrometer (HEPS -- particles greater than 30 keV) were declared
successful on UARS day 2670 (January 2, 1999). Beginning this date, all PEM
activations should yield correct scientific data; however, data generated from
the HEPS subsystem between December 17, 1998 and January 1, 1999 should be
considered invalid and not used. The PEM team would like to thank the FOT, and
in particular Peter Thomas, for their work and efforts related to obtaining a
correct OBC activation procedure for the HEPS subsystem.
***********
HALOE
***********
HALOE was turned on at the beginning of day today. All engineering
parameters are nominal.
HALOE Data Version 19 reprocessing is complete and after the 1997 data is
validated (end of January)the complete data set will be ready for
certification.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
HRDI experienced a flight software error, possibly the result of an upset,
at 01:21 UTC on 7 Jan 99. Data collection was interrupted in response to
the error until 00:00 UTC 8 Jan 99. When data collection was resumed, the
tuning coefficients required were not properly initialized. Recovery from
this error required a reboot and reload which was accomplished on 12 Jan
99. Normal operations resumed at 00:00 UTC 13 Jan 99. Data collected
between 7 Jan 99 and 13 Jan 99 are not valid.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 457 CD-ROM requests as of 15-January-1999.
There was one order received and shipped. There are no longer any remaining
CD-ROM volumes. Should anyone else send a request to receive UARS CD-ROMS,
they will be given instructions as to how to log onto the DAAC for the data.
DAAC transfer data up to October 1998, should be visible to users through the
DAAC interface.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
151 days SUS (7.00) PROD (L3) (2540R-1426C-7F=1107L)
The CDHF upgraded CMS to version 3.9-2 on CDHF2 as part of the
ongoing Y2K implementation phase. Installation Verification Procedure
(IVP) for Decset Environment Manager completed successfully.
CDHF pending software upgrades: IDL 5.03, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4.
***********
FOT
***********
This report covers 02 January 1999 (Orbit 39941, GMT Day 002) through 08
January.
SPACECRAFT OPERATION
The UARS spacecraft passed a milestone of 40,000 operating orbits on 05
January 1999.
The observatory is now in Forward Flight (Southern Hemisphere viewing) with
a +0.5 degree Roll (cold side down) and is performing nominally under two-
battery operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 006/0544 40005 On
HALOE 006/2309 40016 Off
The beta angle ranged between 29.9 degrees to 11.6 degrees for this report
period. The beta angle is now increasing toward a minimum angle of +1.0
degrees on 11 January 1999 (DOY 011).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.6 degrees C. Additional battery performance data
for each battery on 01 January 1999 and 08 January 1999 are provided in
Appendix D.
The clock error ranged from 11.0 to -10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HALOE 01
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/362 - 2287.496912
Transponder B: 99/357 - 2287.499093
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/007 show the observatory data loss to be
21 hours, 41 minutes, 1.332 seconds (an increase of 2.368 seconds since
last report period). This is a 0.0339 percent data loss which equals a
99.9661 percent data capture for the mission.
The increase of 2.368 seconds occurred over five (5) days:
7 0.544 seconds on Day 003 due to hit in data (IP);
7 0.244 seconds on Day 004 due to data hit;
7 0.768 seconds on Day 005 due to no overlap;
7 0.608 seconds on Day 006 due to hit in data (IP); and
7 0.224 seconds on Day 007 due to data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The return to normal attitude (Roll 0.5 to 0.0 degrees) is scheduled for
execution at 011/13:12:11 UTC shortly before the Yaw Around Maneuver.
The next UARS Yaw Around Maneuver is scheduled for scheduled for 11 January
1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 January 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39942 002/0036 99-005 OBC Clk Adj fr 0.00 to 20.6 msec/day
39965 003/1233 99-006 OBC Clk Adj fr 20.6 to 0.00 msec/day
40005 006/0544 99-007 WINDII Power ON
40016 006/2309 99-008 HALOE Power OFF
40043 008/1833 99-009 WINDII WINDII HALT & RESUME
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 January 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI PTS Current Monitor 39964
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
(none this report period)
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 January 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39942 002/00:44 (02 Jan 99) from 00.0 to 20.6 msec/day
39965 003/12:37 (03 Jan 99) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 January 1999
01 Jan 1999 - GMT Day 001 Beta = 29.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +56.0/+33.6 -2.73/-3.99 +0.0/ +0.0 16.8 0.000 0.0
2 +33.6/-22.4 +6.44/+5.15 +31.2/-15.2 25.1 1.026 21.8
3 +44.8/-5.6 +5.15/+3.87 +30.4/-16.0 25.1 1.025 21.6
08 Jan 1998 - GMT Day 008 Beta = 11.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +50.4/+22.4 -3.68/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +84.0/0.0 +5.15/+4.51 +31.6/-14.0 25.6 1.020 21.1
3 +44.8/-5.6 +3.87/+2.92 +30.4/-14.4 25.6 1.018 20.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 January 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-001 FOT 39963 003/0901 14170 Generic Late Acquisition
99-001 FOT 40008 006/0919 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT DEC 98 14170 GENERIC LATE ACQ
98-123 FOT 39788 356/17:45 AP/ODN PROBLEM
98-124 FOT 39878 362/18:00 20485 N0 ODM/GCMR
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-001 FOT see list Jan 98 14170 Generic Late Acq
SEVNTFW99/02